OpenStudy (anonymous):

The coefficient of variation for a sample values Y1,...,Yn is defined by CV=S/Y where Y is the mean and S is the standard deviation. Let Y1,...,Y8 denote a random sample of size 8 from a normal disttribution with mean 0 and variance (sigma)^2

OpenStudy (anonymous):

a) Obtain the distribution of\[8Y/S ^{2}\] b) What is the distribution of \[S^{2}/ 8Y^{2}\] c) Find the number c and such that c>0,d>0 \[P(c \le S/Y \le d) = 0.95\] d) Find the number c such that c>0 \[P(-c \le S/Y \le c) = 0.95\]